A survey of 280 homeless persons showed that 63 were veterans. Construct a​ 90% confidence interval for the proportion of homeless persons who are veterans.
Answer by passionn(5)

z=1.645
p=63/280=0.225
CI=p-z(√(p(1-p) )/n))≤Π CI=0.225-1645(√(0.225(1-0.225) )/280))≤Π<0.225+1.645(√(0.225(1-0.225)/280))
CI=0.184<Π<0.266